# Floquet Schrieffer-Wolff transform based on Sylvester equations

### v2

@@ Abstract
We present a Floquet Schrieffer Wolff transform (FSWT) to obtain effective Floquet Hamiltonians and micro-motion operators of periodically driven many-body systems for any non-resonant driving frequency. The FSWT perturbatively eliminates the oscillatory components in the driven Hamiltonian by solving operator-valued Sylvester equations. We show how to solve these Sylvester equations without knowledge of the eigenstates of the undriven many-body system, using the driven Hubbard model as an example. In the limit of high driving frequencies, these solutions reduce to the well-known high-frequency limit of the Floquet-Magnus expansion. We anticipate this method will be useful for describing multi-orbital and long-range interacting systems driven in-gap.
@@ I Introduction
External driving has become an irreplaceable tool for the control of many-body dynamics in cold atoms [1, 2], correlated materials [3, 4], multi-band electrons [5, 6, 7], photons in wave-guides [8], and hybrid quantum systems [9, 10]. Seminal experimental observations in cold atoms include, inter alia, light-induced gauge fields [11, 12] and topology [13, 14], Floquet-induced many-body localisation [15, 16], discrete time crystals [17, 18], driving-induced superfluid-Mott transitions [19], tuning of exchange interactions [20] and occupation-dependent tunneling [21] in cold atoms. In driven correlated solids, fascinating effects such as transient superconductivity [22, 23, 24, 25, 26, 27, 28], dressed surface states in topological insulators [29, 30], or the light-induced anomalous Hall effect [31] were shown to emerge. Floquet topological insulators can be engineered in acoustic [32] and photonic [33] systems.
@@
In this work, we develop a Floquet Schrieffer Wolff transform (FSWT) which can be applied beyond the HFE. This method constructs an effective Floquet Hamiltonian which is generated perturbatively in the driving strength $g$ rather than the inverse driving frequency. It is, therefore, particularly well suited to describe low-frequency or in-gap driving of multi-orbital systems, as illustrated in Fig. 1. The FSWT derived in this work is conceptually equivalent to a time-dependent transform previously used in circuit QED systems [51, 52], which perturbatively eliminates the time-dependence in the two-body Hamiltonian of the driven Jaynes-Cummings model. Here, we use the derivative of the exponential map to formulate this transform as a systematic many-body perturbation theory based on Sylvester equations, which we can solve with well-controlled approximations for driven many-body systems without knowledge of their eigenbasis. We illustrate our method with the example of a driven Hubbard chain. We demonstrate the importance of correlated hopping terms, whose strength is underestimated in HFE. They can substantially change the Floquet-induced metal-insulator transition when the driving frequency is not too large. Using infinite DMRG simulations and exact diagonalisation, we benchmark the FSWT method and find substantial improvement compared to HFE Floquet approaches in large regions of parameter space (corresponding to the right part of Fig. 1).
@@
Figure 1: Sketch of the application regimes of different Floquet theories in terms of the driven system’s interaction strength and the driving frequency relative to a system resonance: Our lab-frame FSWT, marked by (a), provides a unified result that converges for all non-resonant driving regimes. The high-frequency expansion encounters convergence issues when the oscillation of the Hamiltonian becomes slow. The rotating-frame HFE in Ref. [44], labelled (b), diverges when the correlations are weak. The lab-frame HFE, labelled (c), diverges under the in-gap drive and converges slowly in the strongly correlated case when the interaction is comparable to the driving frequency.
@@ II.1 General Formalism
where $\hat{F}_{t}^{(n)}\sim g^{n}$. We require the unitary $\hat{U}_{t}$ generating the Floquet micro-motion to ($\bf{i}$) preserve the undriven dynamics and ($\bf{ii}$) preserve the macro-motion. Thus, we set
@@
Now, we can determine the functions $\hat{F}_{t}^{(n)}$ for each order $n=1,2,3,...$, which eliminates the time-dependence at order $g^{n}$ in the transformed Hamiltonian $\hat{H}^{\prime}$. This is achieved iteratively: the time-requirement $\partial_{t}\hat{H}^{\prime(n)}_{t}=0$ determines $\hat{F}_{t}^{(n)}$, and $\hat{F}_{t}^{(n)}$ in turn determines the static part of $\hat{H}^{\prime(n+1)}_{t}$ 111 To detail this iteration, we start at $n=1$ in Eq. (9): The requirement $\partial_{t}\hat{H}^{\prime(1)}_{t}=0$ on (9b) eliminates the time-dependence to order $g^{1}$ in $\hat{H}^{\prime}$, from which $\hat{F}_{t}^{(1)}$ is determined. Inserting $\hat{F}_{t}^{(1)}$ back to Eq. (9) not only renders the entire $\hat{H}^{\prime(1)}_{t}$ static in (9b), but also completely determines the static part of $\hat{H}^{\prime(2)}_{t}$ in (9c). Then we proceed to $n=2$: Using the function $\hat{F}_{t}^{(1)}$ derived above, the equation $\partial_{t}\hat{H}^{\prime(2)}_{t}=0$ for (9c) eliminates the time-dependence to order $g^{2}$ in $\hat{H}^{\prime}$, determining $\hat{F}_{t}^{(2)}$. This elimination procedure can be repeated iteratively. Once the $n$-th order equation in Eq. (9) is solved, the function $\hat{F}_{t}^{(n)}$ will be determined which fixes the static part of $\hat{H}^{\prime(n+1)}_{t}$..
@@ II.2 Time-periodic driving
The solution $\hat{f}_{j}^{(1)}$ for all $j$ determines the static part in $\hat{H}^{\prime(2)}_{t}$. This means that, based on the unitary transform $\hat{U}_{t}=\exp\big{(}\hat{F}^{(1)}_{t}\big{)}=\exp\big{(}\sum_{j=1}^{\infty}\hat{f}^{(1)}_{j}e^{ij\omega t}-H.c.\big{)}$, we find a transformed Hamiltonian $\hat{H}^{\prime}_{t}$ with leading oscillatory contribution $\mathcal{O}(g^{2})$. Based on the ansatz (7), we can directly pick out the static part in the transformed Hamiltonian (9) up to order $g^{2}$, which gives
@@
which eliminates the $\mathcal{O}(g^{2})$ terms oscillating at $e^{ij\omega t}$ in $\hat{H}^{\prime(2)}_{t}$. These are again Sylvester equations, determining $\hat{f}_{j}^{(2)}$ and thus the static part of $\hat{H}^{\prime(3)}_{t}$. Collecting the static terms in Eq. (9), we find this static part in $\mathcal{O}(g^{3})$ to be
@@ IV The Driven single-band Hubbard model
As an example, in Fig. 2 (see Appendix E.1 for details), we carry out infinite DMRG simulations of the ground state occupancy, $\langle\hat{n}\rangle=\lim_{L\to\infty}\langle\hat{N}\rangle/L$, in a one-dimensional Hubbard chain using the TeNPy library [71]. In these simulations, we assume a non-zero chemical potential $\mu$. This is necessary to observe the equilibrium metal-insulator transition in one-dimensional Fermi Hubbard model [72], and it does not alter the solution of any of the Sylvester equations. We rescale the horizontal axis to $\mu-U/2$ to match the convention of Lieb-Wu equation. Near criticality, where the charge excitation gap closes and the DMRG algorithm no longer converges, we extrapolate the DMRG result using the Lieb-Wu equation [72]. Our simulations show how the correlated hopping term in our FSWT result Eq. (29), ignored in the lowest order HFE, modifies the occupancy near the Mott transition. We see in Fig. 2 that it weakens the Mott insulating phase. It changes the phase boundary as well as the charge susceptibility of the metallic phase (i.e. the slope of the lines) close to the boundary. The change is comparable in size to the dynamical localisation effect itself, i.e. the shift from the undriven case to the HFE result. We expect this modification to be observable in electronic platforms, for instance, in quantum dot Fermi-Hubbard simulators [58] where $\mu$ is controlled by gating.
@@ IV.3 Comparison to exact dynamics
Figure 3: (a) Return rate of CDW state under driving, Eq. (30), in an $L=10$ site lattice at driving frequency ${\omega}=16J$ and repulsion $U=3J$, with driving strength fixed at $g={\omega}/4$. The Floquet Hamiltonian dynamics are given by TDVP, and the exact time-dependent Hamiltonian dynamics is given by TEBD, with bond-dimension $\chi=300$. (b) Same as (a) but with increased repulsion $U=8J$. (c) The RMSE of the return rate normalised by its mean value, Eq. (31), for $U=3J$, $g=\omega/4$, $L=6$ lattice sites, and a simulation time $tJ=60$.
@@ IV.4 Comparison to Floquet methods in the strongly correlated regime
$\begin{split}&\mathcal{P}(\hat{h}+\hat{H}^{\prime(2)})\frac{1}{-U}(\hat{h}+\hat{H}^{\prime(2)})\mathcal{P}\\ &\approx\frac{1}{-U}\mathcal{P}\hat{h}^{2}\mathcal{P}+\frac{1}{-U}\mathcal{P}\hat{h}\hat{H}^{\prime(2)}\mathcal{P}+\frac{1}{-U}\mathcal{P}\hat{H}^{\prime(2)}\hat{h}\mathcal{P}+\mathcal{O}(g^{4})\\ &=\frac{4J^{2}}{U}(1-2\frac{g^{2}}{{\omega}^{2}})\sum\limits_{j=1}^{L-1}{\bf S}_{j}\cdot{\bf S}_{j+1}\\ &+4\frac{g^{2}J^{2}}{{\omega}^{2}}\left(\frac{1}{U-{\omega}}+\frac{1}{{\omega}+U}\right)\sum\limits_{j=1}^{L-1}{\bf S}_{j}\cdot{\bf S}_{j+1},\end{split}$ (32)
@@ IV.5 Applicability of our model
We have constructed a Floquet Schrieffer Wolff transform to perturbatively generate the many-body Floquet Hamiltonian and micro-motion operators in orders of the driving strength $g$. This transform is obtained from the solution of operator-valued Sylvester equations and does not require knowledge of the eigenbasis of the undriven many-body system. As an example, we showed how to solve the Sylvester equations perturbatively in driven Hubbard models: By expanding the Sylvester equation in orders of hopping $J$, we derived a Floquet Hamiltonian that may be applied for any interaction strength of $U$. This remains true even when $U\sim J$ or $U>{\omega}$, provided that $J\ll{\omega}$ and one does not resonantly drive the Mott gap. We used this Floquet Hamiltonian to investigate how driving-induced interactions modify the Mott-insulator transition in a one-dimensional chain. We showed that these effects, which are underestimated in high-frequency expansions, give a large contribution that can be on par with the dominant renormalization of electronic hopping.
@@
We envision several extensions of our formalism that can treat important scenarios of driven many-body physics: For near-resonant driving, it will be convenient to construct the FSWT in the co-rotating frame, where the solution to the Sylvester equation has no divergence. The corresponding Floquet Hamiltonian then describes how off-resonant driving terms, which give rise to, e.g., the Bloch-Siegert shift [84] in TMDC, influence the resonant dynamics. For deep strong driving (relevant for rf-driven superconducting qubits [85]), where $g\gg{\omega}$, one can apply our FSWT in a nonuniformly rotating frame (see e.g. Refs. [85, 86, 87]), where the time dependence is transferred from $g$ to smaller parameters in the undriven system, e.g. the hopping $J$ (we assume the driving term commutes with the interaction): Constructing the FSWT expansion in powers of $J$, one can describe how the electron hopping perturbatively modifies the strong driving physics.
@@
The FSWT expansion method presented here could be directly adapted to other driven many-body systems with more interaction terms (such as longer-ranged density-density repulsion or multi-orbital Kanamori interactions [88]). This method is also suitable for systems driven by time-dependent interactions [89, 21], where the Floquet Hamiltonian can be obtained by solving Eq. (53). The crucial advantage of our method is to provide a single formula which is not restricted to the driving frequency being the largest energy scale involved. This advantage becomes more prominent when more and more interacting terms are taken into account. Finally, we note that the FSWT provides a convenient and accurate method to analyze driven correlated materials near the Mott transition. This regime is particularly relevant for several unconventional superconductors, where light-induced superconductivity was observed in recent experiments in fullerides [23, 24, 25] and charge transfer salts [27, 28].

### v3

@@ Abstract
We present a Floquet Schrieffer Wolff transform (FSWT) to obtain effective Floquet Hamiltonians and micro-motion operators of periodically driven many-body systems for any non-resonant driving frequency. The FSWT perturbatively eliminates the oscillatory components in the driven Hamiltonian by solving operator-valued Sylvester equations with systematic approximations. It goes beyond various high-frequency expansion methods commonly used in Floquet theory, as we demonstrate with the example of the driven Fermi-Hubbard model. In the limit of high driving frequencies, the FSWT Hamiltonian reduces to the widely used Floquet-Magnus result. We anticipate this method will be useful for designing Rydberg multi-qubit gates, controlling correlated hopping in quantum simulations in optical lattices, and describing multi-orbital and long-range interacting systems driven in-gap.
@@ I Introduction
External driving has become an irreplaceable tool for quantum information processing [1] and quantum simulations [2], as it allows to modify qubit interactions in platforms ranging from Rydberg atoms to Josephson junctions. In addition, it has also become invaluable for the control of many-body dynamics in cold atoms [3, 4], correlated materials [5, 6], multi-band electrons [7, 8, 9], photonic wave-guides [10], and hybrid quantum systems [11, 12]. Seminal experimental observations in cold atoms include, inter alia, light-induced gauge fields [13, 14] and topology [15, 16], Floquet-induced many-body localisation [17, 18], discrete time crystals [19, 20], driving-induced superfluid-Mott transitions [21], tuning of exchange interactions [22] and occupation-dependent tunneling [23] in cold atoms. In driven correlated solids, fascinating effects such as transient superconductivity [24, 25, 26, 27, 28, 29, 30], dressed surface states in topological insulators [31, 32], or the light-induced anomalous Hall effect [33] were shown to emerge. Floquet topological insulators can be engineered in acoustic [34] and photonic [35] systems.
@@
In this work, we develop a Floquet Schrieffer Wolff transform (FSWT) which can be applied beyond the HFE. This method constructs an effective Floquet Hamiltonian which is generated perturbatively in the driving strength $g$ rather than the inverse driving frequency. It is, therefore, particularly well suited to describe low-frequency or in-gap driving of multi-orbital systems, as illustrated in Fig. 1. The FSWT derived in this work is conceptually equivalent to a time-dependent transform previously used in circuit QED systems [53, 54], which perturbatively eliminates the time-dependence in the two-body Hamiltonian of the driven Jaynes-Cummings model. Here, we use the derivative of the exponential map to formulate this transform as a systematic many-body perturbation theory based on Sylvester equations, which we can solve with systematic approximations for driven many-body systems without knowledge of their eigenbasis. We illustrate our method with the example of a driven Hubbard chain. We demonstrate the importance of correlated hopping terms, whose strength is underestimated in HFE. They can substantially change the Floquet-induced metal-insulator transition when the driving frequency is not too large. Using infinite density matrix renormalization group (DMRG) simulations [55] and exact diagonalisation, we benchmark the FSWT method and find substantial improvement compared to HFE Floquet approaches in large regions of parameter space (corresponding to the right part of Fig. 1).
@@
Figure 1: Sketch of the application regimes of different Floquet theories in terms of the driven system’s interaction strength and the driving frequency relative to a system resonance: Our lab-frame FSWT, marked by (a), provides a unified result that converges for all non-resonant driving regimes. The high-frequency expansion encounters convergence issues when the oscillation of the Hamiltonian becomes slow. The rotating-frame HFE in Ref. [46], labelled (b), diverges when the correlations are weak. The lab-frame HFE, labelled (c), diverges under the in-gap drive and converges slowly in the strongly correlated case when the interaction is comparable to the driving frequency. Dashed lines contain the regimes where the HFE methods have better convergence. Note, however, that HFE may still be valid outside these regimes, provided that the driving frequency is the largest energy scale and higher HFE orders are included.
@@ II.1 General Formalism
where $\hat{F}_{t}^{(n)}\sim g^{n}$. We require the unitary $\hat{U}_{t}$ generating the Floquet micro-motion to ($\bf{i}$) preserve the undriven dynamics and ($\bf{ii}$) preserve the macro-motion 111 In the evolution operator described by Eq. (2), the middle part, $e^{-i\hat{H}^{\prime}(t-t_{0})}$, will be identified in Floquet theory as the stroboscopic dynamics, i.e., the Floquet macro-motion. . Thus, we set
@@
Now, we can determine the functions $\hat{F}_{t}^{(n)}$ for each order $n=1,2,3,...$, which eliminates the time-dependence at order $g^{n}$ in the transformed Hamiltonian $\hat{H}^{\prime}$. This is achieved iteratively: the time-requirement $\partial_{t}\hat{H}^{\prime(n)}_{t}=0$ determines $\hat{F}_{t}^{(n)}$, and $\hat{F}_{t}^{(n)}$ in turn determines the static part of $\hat{H}^{\prime(n+1)}_{t}$ 222 To detail this iteration, we start at $n=1$ in Eq. (9): The requirement $\partial_{t}\hat{H}^{\prime(1)}_{t}=0$ on (9b) eliminates the time-dependence to order $g^{1}$ in $\hat{H}^{\prime}$, from which $\hat{F}_{t}^{(1)}$ is determined. Inserting $\hat{F}_{t}^{(1)}$ back to Eq. (9) not only renders the entire $\hat{H}^{\prime(1)}_{t}$ static in (9b), but also completely determines the static part of $\hat{H}^{\prime(2)}_{t}$ in (9c). Then we proceed to $n=2$: Using the function $\hat{F}_{t}^{(1)}$ derived above, the equation $\partial_{t}\hat{H}^{\prime(2)}_{t}=0$ for (9c) eliminates the time-dependence to order $g^{2}$ in $\hat{H}^{\prime}$, determining $\hat{F}_{t}^{(2)}$. This elimination procedure can be repeated iteratively. Once the $n$-th order equation in Eq. (9) is solved, the function $\hat{F}_{t}^{(n)}$ will be determined which fixes the static part of $\hat{H}^{\prime(n+1)}_{t}$..
@@ II.2 Time-periodic driving
Eq. (12) decides the solution of $\hat{f}_{j}^{(1)}$ for all $j$. Based on the unitary transform $\hat{U}_{t}=\exp\big{(}\hat{F}^{(1)}_{t}\big{)}=\exp\big{(}\sum_{j=1}^{\infty}\hat{f}^{(1)}_{j}e^{ij\omega t}-H.c.\big{)}$, we find a transformed Hamiltonian $\hat{H}^{\prime}_{t}$ with leading oscillatory contribution $\mathcal{O}(g^{2})$. Based on the ansatz (7), we know the solution $\hat{f}_{j}^{(1)}$ completely determines the static part in $\hat{H}^{\prime(2)}_{t}$ in Eq. (9c). Thus, we can directly pick out the static part in the transformed Hamiltonian (9) up to order $g^{2}$, which gives
@@
which eliminates the $\mathcal{O}(g^{2})$ terms oscillating at $e^{ij\omega t}$ in $\hat{H}^{\prime(2)}_{t}$. These are again Sylvester equations, determining $\hat{f}_{j}^{(2)}$ and thus the static part of $\hat{H}^{\prime(3)}_{t}$. Collecting the static terms in Eq. (9d), we find this static part in $\mathcal{O}(g^{3})$ to be
@@ IV The Driven single-band Hubbard model
As an example, in Fig. 2 (see Appendix E.1 for details), we carry out infinite DMRG simulations of the ground state occupancy, $\langle\hat{n}\rangle=\lim_{L\to\infty}\langle\hat{N}\rangle/L$, in a one-dimensional Hubbard chain using the TeNPy library [75]. In these simulations, we assume a non-zero chemical potential $\mu$. This is necessary to observe the equilibrium metal-insulator transition in one-dimensional Fermi Hubbard model [76], and it does not alter the solution of any of the Sylvester equations. We rescale the horizontal axis to $\mu-U/2$ to match the convention of Lieb-Wu equation. Near criticality, where the charge excitation gap closes and the DMRG algorithm no longer converges, we extrapolate the DMRG result using the Lieb-Wu equation [76]. Our simulations show how the correlated hopping term in our FSWT result Eq. (29), ignored in the lowest order HFE, modifies the occupancy near the Mott transition. We see in Fig. 2 that it weakens the Mott insulating phase. It changes the phase boundary as well as the charge susceptibility of the metallic phase (i.e., the slope of the lines) close to the boundary. The change is comparable in size to the dynamical localisation effect itself, i.e., the shift from the undriven case to the HFE result. This modification, arising solely from the $\mathcal{O}(J)$ and $\mathcal{O}(J^{2})$ correlated hopping terms, is expected to be observable in electronic platforms, for instance, in quantum dot Fermi-Hubbard simulators [62] where $\mu$ is controlled by gating.
@@ IV.3 Comparison to exact dynamics
Figure 3: (a) Return rate of CDW state under driving, Eq. (30), in an $L=10$ site lattice at driving frequency ${\omega}=16J$ and repulsion $U=3J$, with driving strength fixed at $g={\omega}/4$. The Floquet Hamiltonian dynamics are given by the time-dependent Variational principle (TDVP) [82, 83], and the exact time-dependent Hamiltonian dynamics are given by time-evolving block-decimation (TEBD), with bond-dimension $\chi=300$. (b) Same as (a) but with increased repulsion $U=8J$. (c) The RMSE of the return rate normalised by its mean value, Eq. (31), for $U=3J$, $g=\omega/4$, $L=6$ lattice sites, and a simulation time $tJ=60$.
@@
Several additional benchmarks on the FSWT result are included in Appendices E.2 and E.3: The return rate in the $U>{\omega}$ in-gap driving regime is shown in Fig. 7; the doublon-doublon correlation in the $U\sim{\omega}$ regime is shown in Fig. 8; and the return rate starting from another initial state is shown in Fig. 9. All of our simulations start from product states, as the cold-atom platforms are commonly initialised in such states to study driving-induced dynamics [77, 87]. These simulations further confirm that the FSWT Hamiltonian (29) outperforms the lowest order HFE in the lab-frame, providing a unified description of the non-resonantly driven Hubbard model.
@@ IV.4 Comparison to Floquet methods in the strongly correlated regime
$\begin{split}&\mathcal{P}(\hat{h}+\hat{H}^{\prime(2)})\frac{1}{-U}(\hat{h}+\hat{H}^{\prime(2)})\mathcal{P}\\ &\approx\frac{1}{-U}\mathcal{P}\hat{h}^{2}\mathcal{P}+\frac{1}{-U}\mathcal{P}\hat{h}\hat{H}^{\prime(2)}\mathcal{P}+\frac{1}{-U}\mathcal{P}\hat{H}^{\prime(2)}\hat{h}\mathcal{P}+\mathcal{O}(g^{4})\\ &=\frac{4J^{2}}{U}(1-2\frac{g^{2}}{{\omega}^{2}})\sum\limits_{j=1}^{L-1}{\bf S}_{j}\cdot{\bf S}_{j+1}\\ &~{}~{}+4\frac{g^{2}J^{2}}{{\omega}^{2}}\left(\frac{1}{U-{\omega}}+\frac{1}{{\omega}+U}\right)\sum\limits_{j=1}^{L-1}{\bf S}_{j}\cdot{\bf S}_{j+1},\end{split}$ (32)
@@ IV.5 Applicability of our model
We have constructed a Floquet Schrieffer Wolff transform to perturbatively generate the many-body Floquet Hamiltonian and micro-motion operators in orders of the driving strength $g$. This transform is obtained from the solution of operator-valued Sylvester equations and does not require knowledge of the eigenbasis of the undriven many-body system. As an example, we showed how to solve the Sylvester equations perturbatively in driven Hubbard models: By expanding the Sylvester equation in orders of hopping $J$, we derived a Floquet Hamiltonian that may be applied for any interaction strength of $U$. This remains true even when $U\sim J$ or $U>{\omega}$, provided that $J\ll{\omega}$ and one does not resonantly drive the Mott gap. We used this Floquet Hamiltonian to show that driving-induced correlated hoppings shift the Mott-insulator transition point in a one-dimensional chain. We showed that these effects, which are underestimated in high-frequency expansions, give a large contribution that can be on par with the dominant renormalization of electronic hopping.
@@
We envision several extensions of our formalism that can treat important scenarios of driven many-body physics: For near-resonant driving, it will be convenient to construct the FSWT in the co-rotating frame, where the solution to the Sylvester equation has no divergence. The corresponding Floquet Hamiltonian then describes how off-resonant driving terms, which give rise to, e.g., the Bloch-Siegert shift [91] in TMDC, influence the resonant dynamics. For deep strong driving (relevant for rf-driven superconducting qubits [92]), where $g\gg{\omega}$, one can still apply our FSWT by switching to a nonuniformly rotating frame (via a Peierls transform) [93, 92, 94, 95, 96], where the oscillatory strength is transferred from $g$ to smaller parameters in the undriven system, such as the hopping $J$, see Appendix D.4. In such a frame, we can construct the FSWT expansion in powers of $J$, which provides the stroboscopic Hamiltonian in the $g\gg{\omega}$ case.
@@
The FSWT expansion method presented here could be directly adapted to other driven many-body systems with more interaction terms (such as longer-ranged density-density repulsion or multi-orbital Kanamori interactions [97]). The method is also suitable for systems driven by time-dependent interactions [98, 23], where the Floquet Hamiltonian can be obtained by solving Eq. (53). The crucial advantage of our method is that it provides a single effective Hamiltonian that is not restricted to the driving frequency being the largest energy scale involved. This advantage becomes more prominent when more and more interacting terms are taken into account, such as in driven Rydberg arrays [99, 100] relevant for multi-qubit gate-design [101, 102]. Finally, we note that the FSWT provides a convenient and accurate method to analyze driven correlated materials near the Mott transition. This regime is particularly relevant for several unconventional superconductors, where light-induced superconductivity was observed in recent experiments in fullerides [25, 26, 27] and charge transfer salts [29, 30].